A portion of $6,600 is invested at a 6 percent annual rate of interest, while the remainder is invested at a 4 percent rate. If the annual income from the two portions is the same, what dollar amount was invested at the 6 percent rate?

How can we solve the above question using algebraic equation ? Please help. thanks

Answer by mananth(16946)

6%----------- x
4%-----------6600-x
0.04(6600-x)=0.06x
264-0.04x=0.06x
264 = 0.06x+0.04x
264=0.10x
x=264/0.1
x=2640 Investment @ 6%